Home‎ > ‎Reports‎ > ‎Crew098‎ > ‎


Crew 98 Log Book for January 28, 2011

Photos of the day

Commander's Check-In Report

Virgiliu Pop Reporting

Crew Physical Status: Good; in the Pirate tradition, the Captain got a wood splinter under a nail while singing "Time to say Goodbye"

Time Departed/Returned from EVA: 5 PM-5:20 PM

Brief Narrative of Field Mission Results: Solar clock finished. No EVA report forthcoming.

EVA Data/Interpretations: We put the "5 o'clock" stone on the solar clock.

Engineering/Hab Maintenance: -

Report Transmission Schedule: Intended by 8 PM. Final report of Crew 98 forthcoming tonight.

Plans for Tomorrow: Crew rotation, looking forward to meet the 99ers.

Inventory: We found the missing ATV cover; we had bought a replacement one.

Support Requested:

Miscellaneous: We overused the Internet alloted for today, as i-tunes went, without us knowing, into automatic update mode.

Commander's Report

Virgiliu Pop Reporting

Last full day on Mars, and we are all saddened to leave our alien home. One of us will surely return - Haritina will command the first New Zealand mission to the MDRS, KiwiMars. Last night, she fixed the Celestron - in the tradition of Jeff Hoffman and Story Musgrave, space heroes who fixed the Hubble Space Telescope.

We had a late start today, and the Captain managed to respect the Pirate tradition, getting a splinter under a fingernail while singing "Time to say Goodbye". After Haritina saved Virgil's life (can we bend the truth in our last day here?), four of us went to Hanksville in an attempt to ship some of our stuff to Earth rather than take it home with us on the plane. We all decided for the latter option, as the price for a tiny, 6 pounds box, would have been 300 dollars. About the same price as for sending stuff from the real Mars to Earth!

At our rerturn, we were surprised to see that our computers misbehaved - they connected to i-tunes automatically, exhausting our daily internet quota.

We then started to write our final report, Captain Cook cooked yet another of his famous meals, and here we are, about to compile the final report and send it to the Mars Society headquarters.

Thank you, MS, for two wonderful weeks!

Engineering Report

Haritina Mogosanu Reporting


Kitty (Diesel Generator) (Used? Yes/No): used running all day

Honey (Gas Generator) (Used? Yes/No): not used

Battery Charge Level:

- we have a problem with the voltmeter and it's not working properly on the DC position.

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the

generator on/off):

1 generator run all day ok

Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty):1/2

Propane Tank (Full/0.75/0.5/0.25/Empty):26%

Gasoline Tank (Full/0.75/0.5/0.25/Empty):1 containers 3/4 x 5 gallons


Oil: not checked DG said to leave it on him

Kitty Oil Quantity (# of Quarts in storage):2.5 gal

Honey Oil Quantity (# of Quarts in storage):3 gal

ATV Oil Quantity (# of Quarts in storage):on the engineering work bench

2 x 946 ml

Car Oil Quantity (# of Quarts in storage):where is it?

Water Status:

Outside Potable Water Tank Level (inches from bottom): 7

Trailer Potable Water Tank Level (inches from bottom): 3/4

Hab Potable Water Tank Level (inches from bottom): full

Potable Water Meter Reading: 413995



Condition of plants in Tank 1: plants dead

Condition of plants in Tank 2: duckweed is coming up strongly

Condition of plants in Tank 3: improving

Greenhab notes/ comments:


Telescope Used? (Yes/no) yes

Observatory Notes:



Used (yes/no):Yes

Oil Checked (yes/no):No

Tire Status:not checked

Odometer at end of day:-123527

Notes/Comments on Hab Car:

ATV 1 (Opportunity):

Used: y/n n

Oil Checked: y/n n

Fuel Consumed: filled

Tire Status: not checked


ATV 2 (Spirit):

Used: y/n n

Oil Checked: y/n n

Fuel Consumed:replaced

Tire Status: not checked


ATV 3 (Viking I):

Used: y/n y

Oil Checked: y/n y

Fuel Consumed: replaced

Tire Status: checked


ATV 4 (Viking II):

Used: y/n n

Oil Checked: y/n y

Fuel Consumed: replaced

Tire Status: not checked


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 67

Actual temperature upstairs (Farenheit): 74

Thermostat setting downstairs (Farenheit): 55

Actual temperature downstairs (Farenheit): 64

Computers/Networking Infrastructure:

Notes/comments on internet/computers?: ok - problem with the internet today but will be solved by midnight

General Engineering

1. Notes/Comments:

Cook's Report

Virgiliu Pop Reporting

1a) Today was

- a cooking day, yay!

1b) Was today a special day celebrated at the main meal with special food or activity?

- No.

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

- Morning meal: Apple pancakes, american style, with maple syrup or nutella.

- Lunch: Nothing, the pancakes were very filling.

- Evening meal: Hobo's Caviar

3) List any main meal foods not finished by the crew

- None

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

- Pancakes - just like two days ago, adding some dehydrated apple powder to the mix.

Hobo caviar: Fry a box of Tofu with soy sauce. Hydrate some sliced potatoes, then fry them in a dash of oil on the pancake pan, with red pepper and herbes de provence. Hydrate half a bag (80g, 3 oz) of freeze dried diced ham and some dehydrated leeks, then fry them on the same pan. Last but not least, hydrate some leeks, peas, taco bits (TVP) and dehydrated mushrooms, and fry them together on - guess what? the same pan.

5) Comments/questions for Kim and Jean, the food study investigators

- none